OpenStudy (jamesj):

Logarithms can be surprising. Suppose \( x, y, z > 0 \). Then \[ x^{\log_y z} = z^{\log_y x} \] I'll leave the proof to you.

OpenStudy (bahrom7893):

exercise to the reader ..eh?

OpenStudy (turingtest):

take log_x of both sides

OpenStudy (bahrom7893):

idk why but I'm still uncomfortable with regular logs.. Always have to look up properties

OpenStudy (jamesj):

log base y of both sides. It's not hard. But even though you can write it down in one or two lines, I'm still surprised how this result does not feel intuitive and the kind of thing you'd loose marks for in high school algebra.

OpenStudy (jamesj):

*lose

OpenStudy (turingtest):

I have the proof copied, but I won't post it yet it is easy though...

OpenStudy (mani_jha):

A corollary of this would be: \[x ^{\log_{x}z }=z\] Just take y=z

OpenStudy (mani_jha):

*y=x

OpenStudy (turingtest):

I don't know about log base y, I think I did one ok with log base x of both sides should I post it?

OpenStudy (jamesj):

Como quieras

OpenStudy (turingtest):

nah

OpenStudy (turingtest):

now I just noticed yours is so much simpler anyway James

OpenStudy (jamesj):

For the record then: \[ \huge \log_y( x^{\log_y z}) = (\log_y x)(\log_yz) = \log_y(z^{\log_y x})\]

OpenStudy (jamesj):

the chopped off term is just \[ \huge \log_y(z^{\log_y x}) \]

OpenStudy (mani_jha):

Taking log with base y would give just a 2-line proof \[antilog(\log_{y} z \times \log_{y}x)=antilog(\log_{y}z)^{\log_{x}y } =z ^{\log_{y}x }\]

OpenStudy (anonymous):

\[\Large {x^{\frac{\log_x z}{\log_x y}}} = {z^{\frac{\log_z x}{\log_z y}}}\] \[z - y = x - y \implies z = x\]

OpenStudy (mani_jha):

Hey Ishaan, z=x will make the identity completely meaningless, because then the Left hand side and right hand side will be the one same thing. Are you sure what you did is correct?

OpenStudy (anonymous):

I thought we were supposed to get the solution, my bad :/